### v2

# Tree-Based versus Hybrid Graphical-Textual Model Editors: An Empirical Study of Testing Specifications

Abstract.
Tree-based model editors and hybrid graphical-textual model editors have advantages and limitations when editing domain models. Data is displayed hierarchically in tree-based model editors, whereas hybrid graphical-textual model editors capture high-level domain concepts graphically and low-level domain details textually. We conducted an empirical user study with 22 participants to evaluate the implicit assumption of system modellers that hybrid notations are superior, and to investigate the tradeoffs between the default EMF-based tree model editor and a Sirius/Xtext-based hybrid model editor. The results of the user study indicate that users largely prefer the hybrid editor and are more confident with hybrid notations for understanding the meaning of conditions. Furthermore, we found that the tree editor provided superior performance for analysing ordered lists of model elements, whereas activities requiring the comprehension or modelling of complex conditions were carried out faster through the hybrid editor.
Hybrid Notations, Model Editors, Fuzz Testing, Empirical Study
journalyear: 2024copyright: acmlicensedconference: ACM/IEEE 27th International Conference on Model Driven Engineering Languages and Systems; September 22–27, 2024; Linz, Austriabooktitle: ACM/IEEE 27th International Conference on Model Driven Engineering Languages and Systems (MODELS ’24), September 22–27, 2024, Linz, Austriadoi: 10.1145/3640310.3674102isbn: 979-8-4007-0504-5/24/09ccs: Software and its engineering Domain specific languagesccs: Software and its engineering Visual languagesccs: Software and its engineering Empirical software validation
1. Introduction
Domain-specific languages (DSLs) can have graphical, textual, or hybrid syntaxes. Hybrid graphical-textual DSLs are languages that have a part-graphical and part-textual syntax, i.e., a hybrid graphical-textual syntax. They are particularly appropriate for cases in which one would like to use graphical representations for capturing high-level domain concepts, and textual representations for defining low-level domain details, such as behaviour and complex expressions (Cooper and Kolovos, 2019). Furthermore, they are effectively used via hybrid graphical-textual model editors (Predoaia et al., 2023; Zolotas et al., 2018; Wei et al., 2020), i.e., editors that enable the definition of domain models via graphical and textual representations.
Throughout this paper, we use the term model editor instead of the term workbench; a workbench covers a broader scope, including a model editor and other components like menus, toolbars and perspectives (Burnette, 2005). For brevity, we use the term hybrid instead of hybrid graphical-textual, and the term editor refers to a model editor.
Choosing between a tree-based and a hybrid editor for editing domain models is not a straightforward decision, as both have advantages and drawbacks. Tree editors display domain data hierarchically and in a structured manner by exposing parent-child relationships. Diagrammatic editors contained within hybrid editors do not provide explicit structure, and they are not appropriate for displaying parent-child relationships of multiple levels. For instance, parent elements can be displayed as containers that incorporate their child elements. However, if the children have corresponding child elements, which in turn have child elements as well, displaying all these elements in a diagram can hinder comprehension, as they would effectively be displayed as a container that contains other containers, that in turn contain other containers incorporating a set of leaf child elements. Thus, in such cases, it may be preferred to use multiple diagrams, e.g., by instantiating new diagrams containing the child elements of a parent element. Consequently, the domain data is split across multiple diagrams, and users may face difficulties in being able to relate them. A common drawback of hybrid editors is the initial cost of learning the associated part-graphical and part-textual syntaxes. However, once learned, models may potentially be edited faster, considering also that typically fewer clicks are required due to the definition of textual expressions (Addazi et al., 2017).
Acknowledging the importance of selecting the most suitable editor to support effective and efficient modelling, in this paper we investigate the tradeoffs between tree-based and hybrid editors. System modellers often take for granted the assumption that hybrid notations are superior, however, this claim is not yet supported by empirical evidence. To evaluate this implicit assumption, we conducted an empirical user study to analyse the editor type preferred by system modellers. We recruited 22 subjects, that are either students or professionals working in industry or academia, with various levels of knowledge and expertise in computer programming, modelling, and model-driven engineering (MDE). In the user study, the participants used a DSL that enables the definition of testing specifications of robotic systems in simulation. Each participant used the DSL via both a tree editor and a hybrid editor to execute a set of tasks that were inspired by an industrial case study of the large European Commission research project SESAME (Secure and Safe Multi-Robot Systems) (SESAME, 2024). The user study aims to evaluate the performance and preferences of users, and provide insights on the usage of hybrid editors and tree editors.
Our empirical study revealed that users largely prefer hybrid editors and are more confident with hybrid editors for understanding the meaning of conditions. From a performance perspective, the results pinpoint that hybrid editors are superior for activities involving the comprehension or modelling of complex conditions, whereas tree editors are more efficient for tasks concerning the exploration and analysis of ordered lists of model elements.
Paper structure. Section 2 introduces the necessary background, Section 3 describes our industrial case study, and Section 4 presents the fuzzing-based testing DSL. Sections 5 and 6 detail the methodology and results of the user study, respectively. Section 6.4 identifies threats to validity. Section 7 presents related work and, finally, Section 8 concludes the paper, providing future work directions.
2. Background
Modelling Editors. Hybrid editors are commonly composed of graphical diagram(s), embedded textual editor(s) with syntax-aware editing features, such as syntax highlighting, auto-completion, and refactoring, a view in which one can edit the properties of a selected model element, a view with various symbols that can be dragged and dropped on top of diagrams to create various types of model elements, and a view that consistently reports errors from the model (Cooper and Kolovos, 2019). Textual expressions are able to reference model elements that are defined graphically in diagrams, and consistency between the two is automatically enforced as the model evolves.
Hybrid DSLs and their supporting editors can be developed using Graphite (Predoaia et al., 2023; Predoaia, 2023), a tool that facilitates the systematic engineering of hybrid editors via automatic code generation. Graphite (https://github.com/epsilonlabs/graphite) uses Sirius (Eclipse Foundation, 2024b) for the specification of the graphical part of the syntax, and Xtext (Eclipse Foundation, 2024c) for defining the textual part of the syntax. Alternatively, projectional editors, such as JetBrains MPS (JetBrains, 2024), can be employed for developing such hybrid editors. The research works from (Scheidgen, 2008; Obeo and TypeFox, 2017; Cooper, 2018; Altran, 2022), which are based on hand-written code, introduce techniques for engineering such kind of hybrid editors. However, they do not address consistency enforcement between textual expressions and graphical model elements. Furthermore, except (Cooper, 2018), no techniques have been proposed for uniformly reporting errors. Other streams of work involve using blended modelling (Ciccozzi et al., 2019) and Capella (Eclipse Foundation, 2024a); however, they do not address hybrid DSLs, as justified in (Predoaia et al., 2023).
Testing Specification. Multi-robot systems (MRS) are distributed and interconnected robotic teams capable of collectively performing tasks that are beyond the competency of a single robot (Parker, 2013). Simulation-based testing (SBT) enables investigating the capacity of an MRS to operate dependably using a virtual environment that closely resembles the real deployment environment. Modern robotic frameworks enable realistic simulation of robots, with strong conformance to physical hardware, making the software deployment to real robots easier (Timperley et al., 2018). Compared to field testing, simulation-based testing supports searching large design spaces of components and system configurations, using limited physical hardware resources, thus reducing significantly the overheads of detecting potential violations of safety requirements (Harbin et al., 2021; Gerasimou et al., 2019).
Fuzz testing is an automated software testing technique that involves providing unexpected, invalid or random data as input to a computer program, aiming to find exceptions, crashes and potential memory leaks (Godefroid, 2020). Fuzz testing is considered among the most efficient approaches for finding security vulnerabilities (Liang et al., 2018). In this work, fuzz testing is applied as a simulation-based testing technique to measure the robustness of a robotic mission.
3. Case Study
The use case is an industrial case study in gearbox manufacturing provided by our SESAME industrial partner, KUKA (KUKA, 2024), a company that manufactures industrial robots and factory automation systems. This use case incorporates mobile robots to assemble gearboxes, together with human operators surrounding the assembly cell who assist in the process. During testing, we aim to obtain evidence to validate the safety requirement that robots will not generate hazards to human workers during operation. The primary safety consideration is to ensure that robots do not enter regions that could cause a collision hazard to human operators placed around the perimeter of the assembly cell frame. Under normal circumstances, motion trajectories are pre-planned to ensure conformity to this safety requirement. However, faults in communication, sensing, or motor interlock actuation may lead to robots malfunctioning and entering these forbidden areas. The simulation scenario was supplied for testing by end users, and implemented using a proprietary simulation platform provided by our industrial partner and used as a digital twin (Partners, 2023a).
Figure 1. Case study screenshot from MRS simulation
Figure 2. Fuzzing-based testing DSL metamodel (excerpt)
Figure 3. Meta-classes of condition-based fuzzing (excerpt)
Figure 1 shows the layout of the assembly cell. The primary safety requirement is that the robot grippers are forbidden from entering static safety zones surrounding the cell in order to prevent hazards to human workers. The safety zones are illustrated as blue cuboid regions (Z1, Z2, Z3) on the right-hand side of the external cell frame in Figure 1. If the safety zone surrounding the gripper of the right-hand robot (green cylinder labelled ZR3200) collides with these regions, it will trigger violations of the static safety requirement. An output metric is used to quantify violations of the principal safety requirement, by tracking the occurrence of collisions of the gripper safety zone with any of Z1, Z2 or Z3. Following a safety analysis, we have identified the communication from the controller to the robot arm joints as vulnerable to message disruptions. Therefore, the primary safety requirement will be tested by applying fuzzing, altering the target motion value for robot arm joints with targeted uniform random jitter around the intended point.
We employ a DSL to raise the level of abstraction and make testing specification more comprehensible to robotics engineers. Modelling testing specifications via a DSL tailored to this specific domain is highly desirable, enabling non-expert users to easily test robotics systems (Casalaro et al., 2022; Bruyninckx et al., 2013; Harbin et al., 2023; Wood et al., 2020). Participants of the user study will apply the DSL in the context of this case study, via a tree editor and a hybrid editor, to investigate the following research questions:
  1. RQ1
    (Performance) How do users perform when executing certain tasks with a hybrid editor compared to a tree editor?
  1. RQ2
    (Confidence) How confident are users regarding the correctness and completeness of their solutions to given tasks?
  1. RQ3
    (Preference) Do users prefer hybrid editors or tree editors?
4. Fuzzing-Based Testing DSL
This section presents the fuzzing-based testing DSL developed in the context of the SESAME project, and applied to the case study from Section 3. Through the DSL, users can define testing specifications for an MRS simulation, which controls how fuzzing is applied during testing. Next, we present the abstract syntax and the concrete syntaxes of the DSL and the testing process.
4.1. Abstract Syntax
The abstract syntax (i.e., the metamodel of the DSL) has been defined with the Eclipse Modelling Framework (EMF) through the Ecore metamodelling language. Figure 2 shows an excerpt from the metamodel of the DSL as a class diagram. The core meta-classes are TestingSpace, TestCampaign and Test. A TestingSpace (highlighted) represents the root element of the domain and comprises permissible fuzzing operations (possibleFuzzingOperations) that bind the potential fuzzing space. The testing space also includes performance metrics, which are used to quantify the robotic system performance for safety violations. A TestCampaign specifies and constrains an individual experiment, referencing from the TestingSpace a subset of metrics to use in campaign evaluation, together with a subset of chosen fuzzing operations in includedOperations. The performedTests attribute in TestCampaign is automatically populated during the experiment with dynamically generated and executed Tests.
The Test class represents a test configuration, corresponding to a specific selection of fuzzing operations. During Test execution, instances of MetricInstance are recorded, to quantify the values of particular Metrics under the applied fuzzing.
The FuzzingOperation class is a specific fuzzing operation, i.e., an entity that represents a specific strategy for making runtime modifications to the MRS. The property variableToAffect selects the MRS variable to which fuzzing operations are applied. These variables will be automatically manipulated by the testing platform when fuzzing is active. Subclasses of FuzzingOperation represent distinct fuzzing semantics; e.g., RandomValueFromSetOperation enables randomised changes to structured parameters within a message.
A CampaignResultSet has a name and represents either final or partial intermediate testing results, as specified by the ResultSetStatus enumeration. The results are shown as references to notable tests under results. In an evolutionary experiment, the set of non-dominated tests in the final campaign result set constitutes the Pareto front of optimal metrics obtained during the experiment. The inclusion of a Test in the non-dominated result sets shows that no other tests dominate it with respect to the set of metrics used.
Condition-based fuzzing allows users to activate and deactivate fuzzing based on custom scenario-specific events, incorporating complex logic and spatial or geometric properties. A condition is a logical expression built hierarchically from comparison and logical operators, incorporating state variables from the robotic system. The values of condition variables change dynamically during test execution, with every variable value supplied from a Metric of the corresponding name. Activation of a condition-based fuzzing operation is controlled by its start and end conditions. When the start condition of a fuzzing operation holds, that fuzzing operation will be activated, and will manipulate simulation events. Once activated, the fuzzing operation remains active until its end condition holds.
Condition-based fuzzing experiments aim to discover these conditions automatically, identifying high-level MRS events that lead to failures or safety violations. The primary role of a system tester will be to inspect, parse and understand these auto-generated conditions. However, during debugging and exploratory testing, users may modify existing conditions or create new ones from scratch, to verify the impact of these changes on the behaviour of a test.
Figure 4. Tree Editor — Testing Space
Figure 5. Tree Editor — Activation Condition
Figure 6. Tree Editor — Properties View
Figure 7. Hybrid Editor — Diagram of a Testing Space
The metamodel fragment in Figure 3 shows the meta-classes for condition-based fuzzing. The Condition class comprises an abstract ConditionElement, which is a BasicCondition or CompositeCondition. A BasicCondition represents comparisons of a ConditionVariable with a ConditionExpr, which is either a ConditionVariable or ConditionInteger. The BinaryComparisonOperation is either LESS_THAN, GREATER_THAN, or EQUALS. ConditionVariables contain a reference to an associated metric that supplies its value, and ConditionIntegers represent primitive integers. A CompositeCondition contains two other ConditionElements in the left and right branches, together with a BinaryLogicalOperation. These logical operations permit AND or OR operators. Since condition expressions are dynamically generated during experiments, CompositeCondition permits complex multi-level boolean expressions of arbitrary depth.
4.2. Concrete Syntaxes
The fuzzing-based testing DSL can be used through two concrete syntaxes: a graphical tree-based syntax and a hybrid syntax. The graphical syntax is provided by the default EMF-generated tree editor, whereas the hybrid syntax was developed using Graphite (Predoaia et al., 2023).
Tree editor. A subset of a testing space used in the case study is shown in Figure 4. The TestCampaign is selected in the editor, and its attributes (e.g., includedOperations and metrics) are displayed in the Properties View from Figure 6. The test campaign contains three executed tests and two campaign result sets. Recall that a CampaignResultSet represents a logical grouping of executed tests. The first test contains the RandomValueFromSetOperation under includedOperations in the TestCampaign, and also two metric instances: one for fuzzingOperationTimes and one for collisionOccurrence, as specified by the metrics attribute displayed in the Properties View from Figure 6. Figure 5 expands the RandomValueFromSetOperation named attackJointPosR3200ProductMove_Link0 and displays its child elements. This operation is activated when its start condition holds. The Condition element is the root of the activation condition and it has a CompositeCondition, in which the left side is a BasicCondition and the right side is another CompositeCondition.
Figure 8. Hybrid Editor — a test with a condition textual expression (left diagram) and campaign result sets (right diagram)
Hybrid editor. The Graphite-based concrete syntax is shown in Figures 7 and 8. A graphically defined testing space is illustrated in the diagram in Figure 7. The elements of the testing space are displayed as rectangles with various colours and symbols depending on each element type. Arrows highlight relationships between elements. To create new elements, users can drag and drop various symbols from the Palette on top of the diagram. The diagram displays the high-level elements of the testing space; however, users can view lower-level details by double-clicking on specific elements. By double-clicking on a Test contained in a TestCampaign, a new diagram is instantiated with the child elements of the Test. For instance, Figure 8 (left diagram) shows a hybrid editor that contains the child elements of the first test from the test campaign. The activation condition is specified with a textual expression in an embedded textual editor, with developer assistance features such as syntax highlighting and auto-completion. Users can also double-click the TestCampaign from Figure 7 to view its campaign result sets in a newly instantiated diagram. Figure 8 (right diagram) illustrates the campaign result sets of the test campaign, ordered from first to last via arrows. By double-clicking a Test from the campaign result sets, a new diagram similar to the one from Figure 8 (left diagram) is instantiated, illustrating the details of the Test. We should note that the capabilities and behaviour of the hybrid editor, e.g., the instantiation of new diagrams via double-clicking, are shown to the participants of the user study via a video tutorial.
4.3. Simulation-Based Testing
The SESAME tools for simulation-based testing (Partners, 2022, 2023b) provide a model-driven code generation engine. Input models conforming to the DSL configure an evolutionary optimisation loop for a given TestCampaign. The loop optimises a population of fuzzing tests with the intent of maximising safety requirement violations. For every Test generated and added to the model, the SESAME code generation engine generates a test runner, including an MRS logical interface enabling two-way runtime communication. The test runner launches the simulation, monitors and manipulates fuzzing variables to perform the operations defined for that test. Performance metrics are computed from monitored variables, and their final values are logged under Test as a MetricInstance.
Table 1. Breakdown of expertise in participants
Expertise Modelling DSLs MDE MRS
None / Low 6 7 8 17
Medium 6 8 4 3
High 10 7 10 2
5. Study Methodology
5.1. Recruitment
As the study was part of a wider research project with a broad range of industry and academic partners in modelling platforms, we emailed an invitation letter to target individuals with relevant experience, with a request to indicate their availability for the 2-hour-long study. We shared the same email invite within the Computer Science department of the University of York, UK, for undergraduate and postgraduate students and academics with experience in computer programming, modelling or MDE. The invitation was also shared with various research groups in academia and professionals working in the industry, with modelling expertise.
5.2. Participants
A total of 22 participants participated in the study. The participants currently work either in the industry or academia: 5 participants work in the industry (3 companies across the UK, Germany and Romania), and 17 participants come from academia across 8 universities in the UK, Spain, Austria and Germany. Two of the academic participants were computer science students. Participants had a range of years of programming experience, with 10 participants with more than 10 years, 5 participants with 7–10 years, 5 participants with 4–6 years, and 2 participants with less than 1. Table 1 shows the level of knowledge and expertise of participants.
5.3. Procedures
The study was pilot-tested with two users who did not participate in the final study. During the pilot, we identified practical issues and revised the study accordingly, e.g., to reduce the time needed for the study and improve the wording of task descriptions. Once a participant agreed to join the study, the first author would contact them to arrange a time for the user study. A facilitator, i.e., the first author, guided and observed participants during the sessions. The faculty research committee ethically approved the study.
Each participant joined a virtual meeting with video conferencing together with the facilitator via Zoom. At the beginning of the session, the facilitator described the context of the user study and shared two documents with the participant: an information sheet describing the study and a consent form. The participant read these documents and then signed the consent form if they agreed to continue with the user study session. Next, the facilitator shared a 20-minute video tutorial, presenting the abstract and concrete syntaxes of the fuzzing-based testing DSL and practical examples of its usage in the Eclipse Integrated Development Environment (IDE) via the hybrid and tree editors. After watching the video, the participants could ask questions on the presented information.
Each participant completed a Google Forms survey, which contained the tasks and follow-up questions (e.g., regarding the correctness and confidence in their answers/solutions). Throughout the user study, the facilitator clarified any uncertainties related to the task descriptions. After completing all tasks, participants were asked to report their preferences, including features they liked and disliked about the editors. Lastly, the participants completed a demographic survey, including details of their years of experience and level of expertise in the areas of Table 1.
The participant connected remotely to the computer of the facilitator via TeamViewer. The computer ran a Windows Operating System (OS). On the computer, two Eclipse IDE instances were running for defining and editing testing specifications: one containing the tree editor and another containing the hybrid editor. Participants executed given tasks through both editors using the same semantic model. To reduce memorising a task solution, participants were not told that the underlying semantic model is the same.
To reduce bias and raise the generalisability of the results, the order in which the two types of editors were used for a specific task was interchanged between participants. Half of the participants first carried out a given task with the hybrid editor, and then performed the same task once more with the tree editor. For the other half, the order was reversed, thus participants first executed the task with the tree editor. By alternating the order we can reduce its effects on performance, confidence, or preference, as a task is typically performed more efficiently and faster the second time than the first.
To meticulously analyse how users performed, all sessions were audio- and video-recorded. The participants were encouraged to think aloud while completing the tasks, to record and later analyse any relevant comments. Also, a tool for visually highlighting mouse clicks was used to display a red-coloured circle around the mouse pointer when a mouse click occurred. This allows for counting only relevant mouse clicks when analysing the video recordings. For example, mouse clicks outside the editors are not relevant.
Table 2. Tasks for Understanding (U) and Modelling (M)
ID Description
U1 In the first test, find the Result Value of the Metric Result associated with FuzzingOperationTimesMetric.
U2 In the first test, in the RandomValueFromSetOperation, describe the activation starting condition.
U3 In the campaign result sets of the TestCampaign, find the differences among the last 3 NONDOM sets.
U4 In the testing space, find the variable(s) related to the StreamMetric named collisionOccurrence.
U5 In the testing space, find the fuzzing operation with the largest DoubleRange interval.
M1 In the testing space, create a new TestCampaign and set its name as NewTestCampaign.
M2 Copy the first test from the prior existing TestCampaign into the NewTestCampaign.
M3 In the RandomValueFromSetOperation, modify the activation starting condition by replacing the operators.
M4 In the RandomValueFromSetOperation, change the first referenced variable of the activation starting condition.
M5 Define from scratch the activation ending condition for the RandomValueFromSetOperation.
5.4. Tasks
Participants were given the tasks from Table 2, categorised into two types. Tasks U1U5 focus on understanding a testing space, used to evaluate which type of model editor is better for comprehension. Tasks M1M5 focus on modelling a testing space, used for evaluating which model editor provides better performance when modelling. The table shows a summary of the tasks; the complete task descriptions used in the study are available in (Predoaia, 2024). All given tasks were driven by the industrial case study, aiming to reflect realistic modelling scenarios performed in practice by domain experts while covering a wide range of modelling activities such as finding specific elements, finding relationships and differences between elements, changing references, and creating and copying elements. Participants operate over the semantic model presented in Figure 4, which contains a subset of the relevant model elements from the actual model used in the case study. All tasks of the same type were first executed with one editor, and then they were executed again with the other editor, e.g., after completing all modelling tasks with the hybrid editor, they were executed again with the tree editor.
Understanding Tasks (U1 – U5). To perform the tasks using the tree editor, participants have to carefully inspect all model elements, and accordingly expand them to view their child elements and scan their attributes displayed in the Properties View. We describe next how a user should execute the tasks via the hybrid editor.
Task U1 requires participants to locate various model elements with specific values of interest. More specifically, the task asks participants to locate the first test contained in the test campaign from Figure 7 and view its details through a newly instantiated diagram by double-clicking the test. In the instantiated diagram, that looks like that in Figure 8, participants must locate the Metric Result that contains a FuzzingOperationTimesMetric, and then identify its Result Value. For the diagram from Figure 8, the Result Value is 0.01.
Task U2 asks participants to describe conditions defined through a tree representation and a textual representation. The task entails viewing the details of the first test as presented above, locating the RandomValueFromSetOperation, finding the activation starting condition, and describing it. In the tree editor, the activation starting condition looks as in Figure 5, whereas in the hybrid editor, it looks as the textual expression displayed in the embedded textual editor from Figure 8. When using the tree editor, the participant was asked to construct an equivalent textual expression as pseudocode. Correspondingly, when using the hybrid editor, the participant was asked to find the textual expression, describe its meaning, and categorise conditions as either BasicCondition or CompositeCondition, and identify the left and right branches of CompositeConditions.
Task U3 asks participants to report differences or similarities among an ordered list of model elements. Participants should locate the TestCampaign from Figure 7 and analyse its campaign result sets using a newly instantiated diagram by double-clicking the TestCampaign. In the instantiated diagram, like that in Figure 8, participants must find the last 3 campaign result sets whose names are prefixed with NONDOM, and then notice their differences and similarities. Figure 8 contains a subset of campaign result sets from those used in the user study, which included 15 result sets.
Task U4 asks participants to find relationships (references) between elements. Participants should find the collisionOccurrence StreamMetric from Figure 7, and then identify the related variables. For instance, as indicated by the arrows in Figure 7, the StreamMetric is related to the variables safetyzones/Z1 and safetyzones/Z2.
Task U5 asks participants to find specific types of model elements and perform minor mental computations. Participants should locate the fuzzing operation with the largest DoubleRange interval. Note that the largest interval is that with the highest difference between the upper bound and the lower bound. Although multiple fuzzing operations existed in the testing space of the user study, in Figure 7 there is only one fuzzing operation: the RandomValueFromSetOperation, which has a DoubleRange interval of [-0.3; 0.4].
Modelling Tasks (M1 – M5). In these tasks, participants edit existing and create new model elements, and change relevant property values via both editors. The modelling tasks are done after completing the understanding tasks; thus, participants have a good understanding of the testing space before engaging in these tasks.
Task M1 requests participants to create a new model element and set one of its properties. For instance, in the hybrid editor from Figure 7, the participant would have to identify in the Palette the symbol related to TestCampaign, then drag and drop it on the diagram, and lastly set its name using the Properties View.
Task M2 requires participants to execute operations within the model editors, specifically, the copy and paste operations. The participant would have to either use the context menu or a keyboard shortcut to trigger the copy operation of the first test, and then execute the paste operation into the new test campaign.
Task M3 asks participants to modify an existing condition, through its tree representation as in Figure 5 and its textual representation as in Figure 8. More specifically, the task requires the participant to replace all AND operators with OR operators, and replace the comparison operators LESS_THAN and EQUALS with GREATER_THAN.
Task M4 requests participants to change a reference within a condition. In particular, the task involves changing the first referenced metric from the activation starting condition with another metric. For instance, the first referenced metric in the condition from Figure 8 is collisionOccurrence, and participants had to use fuzzingOperationTimes instead. This change would be realised in the tree editor by changing the reference via a selection menu, whereas in the hybrid editor, this would be done by either typing and deleting text manually, or by using the auto-completion menu that lists all metrics that can be referenced.
Task M5 requires participants to define a condition from scratch that meets certain requirements. Participants should define the activation ending condition such that it contains one root CompositeCondition containing two BasicConditions. When defining conditions as in Figure 5, multiple clicks are required to create new elements, whereas via the hybrid editor users can leverage the auto-completion menu to define the conditions textually.
5.5. Data Collection and Analysis
We analysed the video recordings and the surveys completed by the participants to derive the following metrics:
Duration. The duration quantifies the amount of effort required to do a task, measuring the time-on-task, in seconds, within the editors. The duration does not include time spent reading task descriptions, or time wasted when the remote access and control session via TeamViewer was lost. Participants alternated between the Google Forms survey and the Eclipse IDE instances containing the editors, thus we were able to monitor the time spent on each.
Correctness. After performing all user study sessions, we examined whether participants completed the tasks correctly or made mistakes. Thus, we marked solutions as either 1 (correct and complete) or 0 (mistakes). When a task was not solved correctly, we recorded the specific mistake(s) made.
Clicks and Keystrokes. Mouse clicks and keystrokes reflect modelling effort (Hill, 2011). The number of clicks and keystrokes was recorded for the tasks related to conditions modelling, i.e., tasks M3 and M5.
Confidence. After performing each task, participants self-reported their confidence level in their solution, ranging from 1 (very low confidence) to 5 (very high confidence).
Errors Corrected. We recorded the mistakes made in the context of a modelling task, which were later corrected by the user during the execution of the same task and before its conclusion. A mistake represents a performed action that is incorrect, redundant or certainly unrelated to the solution of a modelling task. Note that the clicks and keystrokes and the time spent while making mistakes and correcting them were included in the analysis.
6. Results and Discussion
6.1. Performance (RQ1)
We evaluate research question RQ1 when carrying out various tasks using the tree-based and hybrid editors by considering the metrics: duration, correctness, errors corrected and clicks and keystrokes.
Duration. Table 3 shows the absolute differences between the durations for completing the tasks via the tree ($T$) and hybrid ($H$) editors. The table is split into tasks for understanding (U) and modelling (M), and the five numeric columns indicate tasks U1U5 and M1M5. The delta ($\Delta$) subtracts the duration when using the hybrid editor from the duration taken with the tree editor, i.e., $Delta=Duration(Tree)-Duration(Hybrid)$. For each task, the delta of the mean and median durations are displayed. A positive value signifies a higher duration with the tree editor compared to the hybrid editor; the opposite holds for a negative value. In addition, Figure 9 shows boxplots with the duration needed for executing the understanding and modelling tasks using both editors.
Figure 9. Boxplots — Durations (seconds) for understanding (U) and modelling (M) tasks using the tree (T) and hybrid (H) editors
Table 3. Deltas of Mean and Median Durations (seconds)
Type $\mathbf{\Delta}$ $(\boldsymbol{T}-\boldsymbol{H})$1 #2 #3 #4 #5
U Mean 1.55 315.95 -15.32 10 36.45
Median 1.50 301 -23 0.50 32.50
M Mean 0.59 -9.50 27.09 8.94 9.51
Median -0.50 2 34 3.50 19.50
Completing task U1 required on average (median) 1.55 (1.50) seconds more with the tree editor than with the hybrid editor. Thus, for solving task U1, which requires locating model elements with specific values of interest, the hybrid editor had slightly better performance; however, no significant difference between the two exists. For task U2, the hybrid editor was faster on average by 315.95 seconds, indicating that using textual representations for expressing conditions improves comprehension significantly. In task U3, the tree editor was faster by 15.32 seconds than the hybrid editor; the median value outlines a higher difference, of 23 seconds. Concretely, the relative difference reported to the mean and median durations of the hybrid editor was 15% and 27%, respectively. This shows that tree editors are more appropriate to explore and analyse ordered lists of model elements, as navigating long ordered element lists in diagrams often incurs overheads. For task U4, the hybrid editor was faster by 10 seconds on average, however, this difference is negated by the delta median value, showing the existence of some outliers for this task. We conclude that no significant difference exists in performance between the editors when executing a task that requires finding references between elements. The hybrid editor yielded faster results in task U5, showing that finding elements that meet certain criteria is quicker in diagrams, as users have to meticulously expand and explore all elements in a tree editor.
Considering the modelling tasks, as indicated by the mean and median delta values of tasks M1 and M2, we cannot conclude that there is a difference in performance between the editors for these tasks. For task M3, the hybrid editor resulted in improved performance, indicating that modifying existing conditions via a textual editor with syntax-aware editing features is quicker than through a tree editor. In task M4, the hybrid editor was faster on average by 8.94 seconds, however, the delta value of the median suggests a negligible difference between the performance of the editors. Thus, for solving a task that requires changing a reference, the editors have similar performance, although the hybrid editor may be slightly faster. For task M5, the hybrid editor was faster on average (median) by 9.51 (19.50) seconds, with a relative difference of 8% (16%), thus the hybrid editor was more effective in modelling conditions.
Finally, we used the non-parametric Wilcoxon signed rank test to perform statistical analysis between the duration taken by participants to complete the understanding and modelling tasks using the editors. Our statistical analysis with ($\alpha$=0.05) demonstrated significant differences in understanding tasks U2 (p=0.00001) and U5 (p=0.0127) and modelling tasks M3 (p=0.0002) and M5 (p=0.0329). These results are also underpinned by the boxplots in Figure 9.
Correctness. All participants solved task U1 correctly through the hybrid editor, whereas three participants solved the task incorrectly with the tree editor. From those, two participants reported an incorrect value as a solution, whereas one participant did not find a solution and abandoned the task. For task U2, two participants reported incorrect solutions with the hybrid editor, not being able to accurately state the left and right branches of the conditions represented textually. Correspondingly, seven participants solved the task incorrectly with the tree editor, as they constructed incorrect equivalent textual representations of the conditions from the tree editor by either grouping sub-conditions incorrectly, missing operators or using unsuitable operators. All participants solved correctly tasks U3U5 using both editors. The results evidence that the hybrid editor surpasses the tree editor in correctly locating specific model elements and in comprehending conditions.
In modelling tasks, one participant incorrectly performed task M1 via the hybrid editor. The participant tried to set the name of a TestCampaign by double-clicking its name on the diagram, and then entering the name in the pop-up menu displayed after double-clicking. The participant forgot a detail shown in the video tutorial: double-clicking specific elements in the hybrid editor instantiates new diagrams. As such, the participant instantiated a new diagram with the name entered in the pop-up menu, without noticing that the name of the TestCampaign had not been set. Correspondingly, all participants solved task M1 correctly through the tree editor.
Task M2 was solved correctly by all participants through both editors. One and six participants solved task M3 incorrectly, with the hybrid editor and tree editor, respectively. The most encountered mistake made through the tree editor was that participants did not expand all condition elements to view their child elements, to be able to locate the operators that needed to be replaced. All participants solved task M4 correctly with the hybrid editor, while only one participant failed using the tree editor. For task M5, two participants incompletely defined the required condition from scratch with the tree editor, while one participant made a mistake when textually defining the required condition; specifically, they used an AND operator instead of an OR operator. Consequently, we conclude that there is no significant difference between the editors for solving the tasks, with the exception of task M3, for which the hybrid editor yielded better results. Fine-grained changes to conditions require a high level of attention to detail, and the results indicate that it is more appropriate to modify existing conditions through a textual representation rather than a tree representation.
Errors Corrected. The errors corrected have been recorded for task M5, in which participants defined conditions from scratch. During the condition definition, 14 and 8 participants made incorrect actions with the hybrid editor and tree editor, respectively; the mistakes were corrected later in both cases. The results highlight that more than half of the participants faced challenges during modelling with the hybrid editor. The reason for the high number of mistakes is that the embedded textual editor from the hybrid editor allows any invalid input, such as typos. For all participants, in total, 8 errors have been corrected with the tree editor and 17 errors have been corrected with the hybrid editor. The mistakes made with the hybrid editor were minor, e.g, participants typed LESS_THAN instead of GREATER_THAN, typed lowercase OR instead of uppercase, and used curly braces instead of parentheses for defining BasicConditions. However, the errors that have been corrected in the tree editor were of a higher magnitude. For example, participants created a BasicCondition as a root element, then continued modelling the conditions, and later realised that the root should have been a CompositeCondition. Next, participants deleted the BasicCondition, and defined once again the conditions from scratch, as the tree editor does not allow replacing a model element.
Clicks and Keystrokes. Table 4 shows statistics for clicks and keystrokes, recorded for tasks M3 and M5, which involve modifying and modelling conditions. The delta subtracts the mean and median number of clicks or keystrokes when using the hybrid editor from the corresponding values of the tree editor. On average, more mouse clicks (24.32) and fewer keystrokes (32.55) were used for solving task M3 with the tree editor. Similarly, more clicks and fewer keystrokes were needed to solve task M5 via the tree editor. The results are not surprising, as hybrid editors require more typing to define textual expressions, which are defined via clicks through a tree editor.
Table 4. Deltas of Mean and Median No. Clicks, Keystrokes
$\mathbf{\Delta}$ $(\boldsymbol{T}-\boldsymbol{H})$ Clicks M3 Keystrokes M3 Clicks M5 Keystrokes M5
Mean 24.32 -32.55 29.70 -53.49
Median 25 -14.50 30 -48
Table 5. Deltas of Mean and Median Confidence Levels
Type $\mathbf{\Delta}$ $(\boldsymbol{T}-\boldsymbol{H})$1 #2 #3 #4 #5
U Mean -0.68 -1.18 -0.14 0.03 -0.05
Median -0.50 -1 0 0 0
M Mean -0.09 0.09 -0.18 -0.34 -0.32
Median 0 0 0 0 0
6.2. Confidence (RQ2)
We answer research question RQ2 concerning the users’ confidence when using both editors to do the specified tasks by analysing their self-reported confidence level, ranging between 1 (very low confidence) to 5 (very high confidence). Table 5 shows statistics of the confidence level of participants about their task solutions. As before, the table is split into tasks for understanding (U) and modelling (M), and the five numeric columns are linked to tasks U1U5 and M1M5. The delta subtracts the mean and median confidence when using the hybrid editor from the mean and median confidence when using the tree editor. For task U1, participants were on average slightly more confident with their solution when using the hybrid editor, signifying users’ confidence for tasks that require understanding a domain and locating various model elements with specific values of interest. Participants were on average more confident by 1.18 points, with their solution to task U2 when using the hybrid editor, signifying their confidence in understanding the meaning of conditions. For the remaining tasks, the results show no significant differences in the confidence of the solutions.
We also analyse the correlation between incorrect solutions and confidence level, aiming to determine how confident participants are with their incorrect solutions. A mean confidence of 3.13 and median confidence of 3 has been reported for the incorrect solutions to tasks U1U5 when using the tree editor. This shows that participants are neutral about the correctness of their partially incorrect solutions. Further, a mean and median confidence of 4.5 has been reported for the incorrect solutions to tasks M1M5 when using the tree editor. Similarly, a mean confidence of 4.67 and a median confidence of 5 has been reported for the incorrect solutions to tasks M1M5 when using the hybrid editor. These results show that participants were confident with their partially incorrect solutions, as they had not noticed any of their mistakes. This finding suggests that the confidence level is related to the ability of users to spot their mistakes. Participants could identify the understanding tasks they could not solve, and that was reflected in their confidence level for each understanding task. However, for the modelling tasks, the participants wrongly assumed that their solutions were correct, leading to a mistaken increased confidence level.
6.3. Preference (RQ3)
We answer research question RQ3 concerning the preferences of users with respect to the two editor types. Before engaging in the study, participants were asked for their preferred editor type for modelling. As reported, 3 participants preferred a tree editor, 11 participants a hybrid editor, and 8 participants stated no preference. After performing the given tasks, the participants were asked again about their preferences. As reported, 2 participants preferred a tree editor, 16 participants (73%) a hybrid editor, and 4 participants had no preference. The results show that one participant changed their preference from a tree editor to a hybrid editor, and half of the participants with no initial preference now prefer a hybrid editor.
Concerning understanding a testing space that does not contain condition expressions, 12 participants reported their preference for a diagrammatic editor, 7 participants for a tree editor, and 3 participants had no preference. To explore and navigate a testing space that does not contain condition expressions, 13 participants preferred a diagrammatic editor, 7 participants a tree editor, and 2 participants had no preference. To find relationships between elements of a testing space, 19 participants preferred a diagrammatic editor, 0 participants a tree editor, and 2 participants had no preference. To understand the meaning of condition expressions, 17 participants preferred a textual representation, 3 participants a tree representation, and 2 participants have no preference. For modelling conditions from scratch, 16 participants preferred a textual representation, 3 participants a tree representation, and 3 participants have no preference. To modify existing conditions, 17 participants preferred a textual representation, 4 participants a tree representation, and 1 participant has no preference. As reported, 20 participants stated that adding a textual syntax for modelling conditions inside the tree editor would improve usability and modelling efficiency; 2 participants did not share this opinion.
The feature of the tree editor that participants particularly liked was its conciseness and hierarchical representation that evidences the order and containment of elements. Participants also noted that the tree editor provided a better sense of control over the model, enabling faster exploration, especially when using the keyboard. However, several participants reported difficulties in understanding and defining complex conditions with the tree editor. Many participants would have liked the tree editor to use graphical symbols and colour-coding to differentiate between element types.
Concerning the hybrid editor, participants particularly liked the flexibility in defining complex conditions and finding relationships between elements. Several participants liked the visual aspect of the hybrid editor, as they could relate colours and symbols to specific element types. However, some participants did not like the need to switch between multiple diagrams and considered that diagrams would not be readable for large models. Also, the textual syntax of conditions was not favoured by all participants, as it contained multiple kinds of brackets for grouping sub-conditions rather than a single kind, and it did not provide any formatting (pretty printing).
6.4. Threats to Validity
Construct Validity. A construct validity threat is that participants may misinterpret task descriptions, although they could eventually understand the descriptions with the help of the session facilitator. If participants begin a task without completely understanding its description, they will need more time to solve it, and they will make more mistakes that they may later correct. Thus, the performance metrics for duration, correctness, and errors corrected are, by design, influenced by the clarity and understandability of task descriptions.
Internal Validity. The main internal validity threat is that the study was conducted by accessing a remote computer via TeamViewer. For tasks M3M5, participants used keyboard shortcuts, e.g., control+space to display an auto-completion menu in the textual editor from the hybrid editor. Some participants faced challenges, either because TeamViewer randomly did not trigger the actions associated with the keyboard shortcuts or because they did not know how to trigger keyboard shortcuts correctly due to their limited knowledge of the Windows OS. In such instances, durations reported with the hybrid editor were higher than expected.
External Validity. An external validity threat is that the tree and hybrid editors have been evaluated with a single concrete syntax. Using a more appealing hybrid notation, with a different graphical representation and with a more intuitive textual syntax, may yield a better performance. Although the tasks covered a broad range of modelling activities, they were uncomplicated, as almost all participants could solve all tasks correctly. Further experiments would be useful to strengthen the generalisability of our findings, involving hybrid editors that are not Sirius/Xtext-based and other kinds of non-EMF or custom EMF-based tree editors.
7. Related Work
DSLs for Robotics. Model driven notations and toolchains have been used in RobotML (Dhouib et al., 2012), BRICS (Bruyninckx et al., 2013), SmartSoft (Schlegel et al., 2009), RoboChart (Miyazawa et al., 2019), and in engineering robotic perception and control  (Ciccozzi et al., 2017; Schlegel et al., 2010). Developing model-driven solutions for the robotics domain is an established area, with a common mechanism for MDE engineering using finite state machines and statecharts (Gascuena et al., 2012; Paraschos et al., 2012; Skubch et al., 2011). MRS interactions rather than single robots are studied in (Cattivera and Casalaro, 2015). The FLYAQ family of graphical DSLs  (Ciccozzi et al., 2016) model the structure and behaviour of multi-robot aerial systems, considering their spatial and situational surroundings and tasks assigned to robots and their behaviour. FLYAQ is extended with a specification language in (Dragule et al., 2017), enabling the definition of domain-specific constraints in a declarative manner and (Pinciroli and Beltrame, 2016) proposes a textual DSL for specifying the behaviour of robot swarms.
Testing DSL Tools. FuzzFactory  (Padhye et al., 2019) is a DSL for developing domain-specific fuzzing applications. It generates code necessary for fuzzing applications to communicate with a fuzzer, and obtain dynamic domain-specific feedback during testing. VerifAI (Dreossi et al., 2019) proposes a complementary approach for tool-supported testing of cyber-physical systems, typically integrated with AI and machine learning approaches. VerifAI uses Scenic (Fremont et al., 2019), which presents a scenario testing DSL, capable of specifying probabilistic features that can be combined to generate test cases of interest.
DSL Evaluation. FQAD (Kahraman and Bilgen, 2015) introduces a methodology for DSL evaluation, considering quality characteristics like expressiveness, usability, and productivity to assess DSL suitability. Despite its usefulness in multiple case studies, FQAD relies on potentially subjective user perspectives, and also lacks empirical validation in textual languages. Similarly, our empirical study relies on subjective data such as user preferences; however, it is also based on objective information, e.g., correctness, number of clicks, errors corrected.
DSL evaluation for multi-agent systems is used in (Challenger et al., 2016), analysing language (elements and transformations), execution and quality. The results show that 80% of the final product code could be auto-generated. However, it did not consider mistakes from language users and their impact on learnability by new users. The textual DSL Athos (Hoffmann et al., 2022) for specifying transport and traffic routing simulations was evaluated to determine its effectiveness and efficiency in model creation and understanding. Athos was found useful in modelling time and test scores when used by users inexperienced in software development, and to a lesser degree, in more experienced users.
User comprehensibility of a business use case between textual and graphical flowchart (BPMN) representations is explored in (Ottensooser et al., 2012). All user groups could understand the textual representation (especially strong readers) and experienced domain analysts obtained increased comprehension from BPMN. This suggests comprehension can be optimised by showing both representations; first textual and then graphical. This may not apply directly to our study where both representations are significantly graphical, with the textual elements restricted to a subset of the syntax (condition expressions).
In (Jakšić et al., 2014), tests were performed to determine whether users could better create and manage feature models (FMs) using a diagrammatic representation, as compared to a textual representation. The results are aligned with our insights and showed statistically that users authored FMs of a higher quality and managed FMs with better efficiency and effectiveness with the diagrammatic representation. Finally, in (Jolak et al., 2020), comparisons were performed between a graphical UML language and textual descriptions of the associated system, a familiar MVC fitness centre architecture. It found that the graphical representation is generally better than a textual description of the associated system, which broadly corroborates the findings in our investigation.
8. Conclusions and Future Work
Our empirical study, using an industrial use case, aimed to analyse and understand users’ performance, confidence, and preferences between using tree-based and hybrid graphical-textual model editors. Concerning performance, there was a tradeoff between the speed of executing tasks using the editors. The hybrid editor was better for comprehending and modelling complex conditions, while the tree editor was better for analysing ordered lists of model elements. When modelling conditions, more corrected errors were found with the hybrid editor; however, the mistakes made were minor. Also, more keystrokes and fewer clicks were made for modelling conditions with the hybrid editor. Users were more confident with their solutions when using the hybrid editor to understand the meaning of conditions. Finally, before the study, 50% of participants preferred a hybrid editor; at the end, the percentage increased to 73%, showing an increased preference towards the hybrid editor.
We plan to validate the generality of our findings by using multiple graphical tree-based and hybrid syntaxes. Also, tree-based and hybrid editors should be evaluated with other domains and case studies, by carrying out a wider range of modelling activities.
Acknowledgements. We thank our SESAME partners and the study participants. This work was supported by the SCHEME InnovateUK project (#10065634), the EU Horizon 2020 project SESAME (#101017258), and the Horizon Europe project SOPRANO (#101120990).